2. Interest rates are currently 2% in the US and 3% in Germany. The current spot rate between the € and $ is $1.5/€. What is the expected spot rate in one year if the international Fisher effect holds? (Answer:1.4854$/€)
Answer #1

SDF 1+1 - 1+i) Itif

S(t+1) = ((1+2%)/(1+3%)) x (1.5)

S(t+1) = 0.9903 x 1.5

S(t+1) = 1.4854